Consider the following two mutually exclusive projects: Year Cash Flow (A) Cash Flow (B) 0 –$356,000 –$40,000 1 31,000 23,000 2 42,000 15,200 3 50,000 14,100 4 445,000 11,200 The required return on these investments is 13 percent. Required: (a) What is the payback period for each project? (Do not round intermediate calculations. Round your answers to 2 decimal places (e.g., 32.16).) Payback period Project A years Project B years (b) What is the NPV for each project? (Do not round intermediate calculations. Round your answers to 2 decimal places (e.g.,32.16).) Net present value Project A $ Project B $ (c) What is the IRR for each project? (Do not round intermediate calculations. Enter your answer as a percentage rounded to 2 decimal places (e.g., 32.16).) Internal rate of return Project A % Project B % (d) What is the profitability index for each project? (Do not round intermediate calculations. Round your answers to 3 decimal places (e.g., 32.161).) Profitability index Project A Project B (e) Based on your answers in (a) through (d), which project will you finally choose?
Answer #1

a

Project A
Year Cash flow stream Cumulative cash flow
0 -356000 -356000
1 31000 -325000
2 42000 -283000
3 50000 -233000
4 445000 212000
Payback period is the time by which undiscounted cashflow cover the intial investment outlay
this is happening between year 3 and 4
therefore by interpolation payback period = 3 + (0-(-233000))/(212000-(-233000))
3.52 Years
Project B
Year Cash flow stream Cumulative cash flow
0 -40000 -40000
1 23000 -17000
2 15200 -1800
3 14100 12300
4 11200 23500
Payback period is the time by which undiscounted cashflow cover the intial investment outlay
this is happening between year 2 and 3
therefore by interpolation payback period = 2 + (0-(-1800))/(12300-(-1800))
2.13 Years

b

Project A
Discount rate 13.000%
Year 0 1 2 3 4
Cash flow stream -356000 31000 42000 50000 445000
Discounting factor 1.000 1.130 1.277 1.443 1.630
Discounted cash flows project -356000.000 27433.628 32892.161 34652.508 272926.834
NPV = Sum of discounted cash flows
NPV Project A = 11905.13
Where
Discounting factor = (1 + discount rate)^(Corresponding period in years)
Discounted Cashflow= Cash flow stream/discounting factor
Project B
Discount rate 13.000%
Year 0 1 2 3 4
Cash flow stream -40000 23000 15200 14100 11200
Discounting factor 1.000 1.130 1.277 1.443 1.630
Discounted cash flows project -40000.000 20353.982 11903.830 9772.007 6869.170
NPV = Sum of discounted cash flows
NPV Project B = 8898.99
Where
Discounting factor = (1 + discount rate)^(Corresponding period in years)
Discounted Cashflow= Cash flow stream/discounting factor

c

Project A
IRR is the rate at which NPV =0
IRR 14.07%
Year 0 1 2 3 4
Cash flow stream -356000.000 31000.000 42000.000 50000.000 445000.000
Discounting factor 1.000 1.141 1.301 1.484 1.693
Discounted cash flows project -356000.000 27176.943 32279.523 33688.892 262854.641
NPV = Sum of discounted cash flows
NPV Project A = 0.000
Where
Discounting factor = (1 + discount rate)^(Corresponding period in years)
Discounted Cashflow= Cash flow stream/discounting factor
Project B
IRR is the rate at which NPV =0
IRR 24.90%
Year 0 1 2 3 4
Cash flow stream -40000.000 23000.000 15200.000 14100.000 11200.000
Discounting factor 1.000 1.249 1.560 1.948 2.433
Discounted cash flows project -40000.000 18415.417 9744.308 7237.361 4602.914
NPV = Sum of discounted cash flows
NPV Project B = 0.001
Where
Discounting factor = (1 + discount rate)^(Corresponding period in years)
Discounted Cashflow= Cash flow stream/discounting factor

d

Project A

PI= (NPV+initial inv.)/initial inv.
=(11905.13+356000)/356000
1.03

Project B

PI= (NPV+initial inv.)/initial inv.
=(8898.99+40000)/40000
1.22

e

Accept project B as it has higher IRR and profitability index